OpenStudy (anonymous):

6 + 2i 3 + 2i Simplify the expression. A) 22 - 6i - 13 B) 14 - 6i - 13 C) 22 - 6i 13 D) 14 + 6i 13

OpenStudy (anonymous):

Use wolframalpha

OpenStudy (anonymous):

google it

OpenStudy (anonymous):

whats wolframalpha

OpenStudy (mertsj):

\[\frac{6+2i}{3+2i}\times\frac{3-2i}{3-2i}=\frac{18-6i-4i^2}{9+4}=\frac{22-6i}{13}\]

OpenStudy (mertsj):

wolframalpha.com where you can ask lots of math questions and get answers to problems.

OpenStudy (anonymous):

is it free?